The Moon has a mass of M = 7.3 · 1022 kg, a radius of R = 1.7 · 106 m and a rotation period of
T = 27.3 days. Scientists are planning to place a satellite around the Moon that always remains
above the same position (geostationary).
(a) Calculate the distance from the Moon’s surface to this satellite.
(b) Explain if such a Moon satellite is possible in reality.
1
Expert's answer
2019-05-13T10:51:26-0400

Gravitational force causes centripetal acceleration


"F = \frac{GMm}{R^2} = m\omega^2 R"

Using "\omega = \frac{2\pi}{T}" , we derive


"\frac{GM}{R^3} = \frac{4\pi^2}{T^2} \Rightarrow R = \sqrt[3]{\frac{GM T^2}{4\pi^2}} = 8.82\cdot 10^7 \,\text{m} = 88200\,\text{km}"

Such a sattelite is impossible in reality, because be couldnt ignore gravitation of Earth.